OpenStudy (anonymous):

Show all Work 1. Solve: y^2 + 7x = 30 2. Solve: x^2 + 7 = -8x 3. Solve: x^2 – 26x + 48 = 0 4. Solve: x^2 - 4x = 32 5. Solve: w(6w + 12) = 0 6. Solve: y^2 + y = 30 7. Solve: (a – 5)(a + 3) = 0

OpenStudy (anonymous):

How to solve them? Is there any method?

OpenStudy (anonymous):

Solving Quadratic Equation

OpenStudy (anonymous):

Ohh okay :) so do you know what's the formula for quadratic equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

LOL okay I'll tell you :) Well listen we have 3 methods of solving quadratic equation

OpenStudy (anonymous):

They're all trivially factorable... I'll do the first one as an example:$$x^2+7x=30\\x^2+7x-30=0\\x^2+10x-3x-30=0\\x(x+1)-3(x+10)=0\\(x-3)(x+10)=0$$Do you know what either \(x-3,x+10\) have to be for their product to be \(0\)? One of them must also then be \(0\) so solve for both cases:$$x-3=0\quad\implies\quad x=3\\x+10=0\quad\implies\quad x=-10$$

OpenStudy (anonymous):

I would not advise teaching him the quadratic formula at this point because these are all solved easily by factoring :-p

OpenStudy (anonymous):

Umm yeahh but the easiest method is by solving quadratic formula :) Well it's his choice :)

OpenStudy (anonymous):

First find the delta: \[\delta = b ^ 2 - 4 * a * c\] After that do to find x' and x'': \[X = -b \pm \sqrt{\delta} \]

OpenStudy (anonymous):

i have no idea of what the equation is

OpenStudy (anonymous):

Solve: x^2 + 7 = -8x x^2+8x+7=0 x^2+1x+7x+7x=0 x(x+1)+7(x+1)=0 (x+1)(x+7)=0 x+1=0 so here x=-1 x+7= so here x=-7

OpenStudy (anonymous):

Umm okay. @AnimalSadqi listen can you understand now we are solving these equations?

OpenStudy (anonymous):

I really would not say it's his choice because in actuality it's his teacher's choice. The problem set strongly suggests he has not learned the quadratic formula yet and instead all lend themselves to factoring (some of them almost *directly*).

OpenStudy (anonymous):

Umm yeahh

OpenStudy (anonymous):

i just need help please!

OpenStudy (anonymous):

Okay sure ask what you want to ask?

OpenStudy (anonymous):

i need help with these questions

OpenStudy (anonymous):

@AnimalSadqi we are solving these equations by breaking the middle term got it?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Suree I'll explain you don't worry :) Just tell me do you know how to break the middle term?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay I'll show you

OpenStudy (anonymous):

See breaking the middle term means that the middle term of the equation like for example we have this equation x^2+8x+7=0 in this equation the middle term is 8x so here we have to break this term(breaking them in such a way that the numbers we are numbers we are taking should plus and make 8x and we multiply both the term they should make 7 Got what I said?

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

 See this drawing and I hope you'll understand :) See here 1+7 makes 8 when we add them but when we multiply them they make 7 Got it?